To fence a rectangular farm of 750m^2, 110m of fence has been used. Calculate the dimensions of the farm.
Answer by mananth(16946)

Area of farm = 750 sq. m
110 meters will be the perimeter
let one side be x
the other side will be 110-2x /2
Area = L * W
x((110-2x /2)=750
x(110-2x /2 ) =750
110x-2x^2/ 2 =750
110x-2x^2= 1500
2x^2-110x+1500=0
x^2-55x+750=0
x^2-30x-25x+750=0
x(x-30)-25(x-30)=0
(x-30)(X-25)=0
x= 30 OR 25
The dimensions of the farm are 35 by 25 meters